# and hierarchical hyperbolicity

### v2

@@ 1. Introduction
Central extensions $1\to K\to E\to G\to 1$ (which we often abbreviate with just $E$) are classified by their associated Euler class $\alpha_{E}\in H^{2}(G;K)$. We call a central extension bounded if its Euler class is bounded, i.e. it is represented by a bounded cocycle. One reason of interest is the fact that bounded central extensions are quasi-isometrically trivial [Gersten_B_is_QIT]. Another reason of interest comes from the study of hierarchically hyperbolic groups, see below.
@@
By a quotient central extension here we mean that there exists a diagram as follows:
@@
By Proposition 2.9 (which is probably known to experts, and was recently obtained independently by Tao and Wan [tao2025properactionsfiniteproducts, Lemma 5.6]), boundedness of a central extension $E$ is equivalent to the existence of a quasihomomorphism $E\to K$ that is the identity on $K$. Quasihomomorphisms, in the sense of e.g. [fujikapo], are also relevant for quotient central extensions, and indeed we show that boundedness of a quotient central extension is equivalent to an extendability problem for a quasihomomorphism. An informal statement is as follows:
@@
We postpone to Section 4 the details of how Theorem 1.3 relates to hierarchical hyperbolicity of quotients of mapping class groups. In the other direction, under a further mild assumption, we prove that a central quotient of a HHG is commensurable to a HHG:
@@
The hypothesis that $E$ has clean containers is mild and natural, as explained in Definition 3.5.
Since $G$ is a finite-index overgroup of a HHG, it inherits the structure of a hierarchically hyperbolic space, given by the quasi-isometric inclusion $E^{\prime}/K\hookrightarrow G$. However, even assuming that $E$ has clean containers, one cannot hope that $G$ itself is a hierarchically hyperbolic group: in Remark 3.15 we argue that the trivial central extension $E=\mathbb{Z}\times G$ of the $(3,3,3)$-triangle group $G$ is HHG, while $G$ is not by [PS_unbounded].
Finally, inspired by the setup explained in Section 4, we show that many quotients of the braid group $B_{4}$ on 4 strands are bounded central extensions. Recall here that $B_{4}$ is a central extension of a finite-index subgroup of the mapping class group of the five-holed sphere. The quotients we consider here are quotients by powers of pseudo-Anosovs, as studied for finite-type surfaces for instance in [DGO, hhs_asdim].
Theorem 1.5 (see Theorem 4.3).
Let $h$ be a pseudo-Anosov element of $B_{4}$. Then there exists $M_{0}\in\mathbb{N}_{>0}$ such that, for all multiples $M$ of $M_{0}$, we have that $B_{4}/\langle\hskip-1.99168pt\langle h^{M}\rangle\hskip-1.99168pt\rangle$ is a bounded central extension of a HHG, and in particular it is a HHG.
@@
Section 3 starts with background material on hierarchically hyperbolic groups (HHG). Roughly, a group $\Gamma$ is a HHG if there exists a collection of hyperbolic spaces $\{\mathcal{C}U\}_{U\in\mathfrak{S}}$, together with coarsely Lipschitz “coordinate projections” $\pi_{U}\colon\Gamma\to\mathcal{C}U$ satisfying various conditions.
@@
We uncover a connection between boundedness and hierarchical hyperbolicity of a central extension. In one direction, Theorem 3.11 shows that a bounded central extension of a HHG is itself a HHG. With an inductive argument, one can reduce to the case where the kernel is $\mathbb{Z}$, which was already settled by [uniform_undistortion, Theorem 5.14]; in turn, the proof of the latter is a refinement of [HRSS_3manifold, Corollary 4.3], which solved the case where the base is hyperbolic. We sketch here the core idea of both arguments, as it enlightens several ideas that appear repeatedly throughout our paper. From a bounded extension one gets an unbounded quasimorphism $\phi\colon E\to\mathbb{Z}$ as above, and the function $E\to\mathbb{Z}\times G$ mapping $e\in E$ to $(\phi(e),\pi(e))$ is a quasi-isometry. In other words, a bounded extension is also quasi-isometrically trivial in the sense of [Gersten_B_is_QIT], though the converse implication is not true in general (see [frigsisto, ascarimilizia] and Remark 3.13). Then the HHG structure for $E$ will have one domain for every domain of $G$, with projection factoring through the quotient map $\pi$, and one quasiline to “detect” the $\mathbb{Z}$-factor. Such quasiline is built out of the quasimorphism $\phi$, using an observation of Abbott, Balasubramanya, and Osin [ABO, Lemma 4.15].
@@
In the opposite direction, Theorem 3.14 proves that, if $E$ is a HHG, then the central extension is bounded, moreover, under the further technical assumption of clean containers, the quotient is a finite-index overgroup of a HHG. To get boundedness we again look for a quasimorphism on $E$ which is the identity on $\mathbb{Z}$. Towards this, one first finds a domain $\mathcal{C}U$ on which the centre acts loxodromically, which must be a quasiline; then the required quasimorphism is the Busemann quasimorphism [Manning_actions_on_hyp, Section 4.1] associated to the action of (a finite-index subgroup of) $E$ on $\mathcal{C}U$, which roughly maps each $e\in E$ to its asymptotic translation length on the quasiline.
@@
We now sketch how to prove hierarchical hyperbolicity of a finite-index subgroup of $G$. Let $\mathcal{C}U_{1},\ldots,\mathcal{C}U_{k}$ be the domains on which $\mathbb{Z}$ acts loxodromically. By taking linear combinations of the associated Busemann quasimorphisms, and then replacing each $\mathcal{C}U_{i}$ with a new quasiline (again using [ABO, Lemma 4.15]), we can modify the HHG structure for $E$ in such a way that $\mathbb{Z}$ acts loxodromically on a single $\mathcal{C}U_{i}$, call it $\mathcal{C}U$, and with uniformly bounded orbits on every other domain. Then the HHG structure of the quotient roughly coincides with what is left of the structure for $E$ after we delete $\mathcal{C}U$.
@@
In the above procedure, one has to restrict to a finite-index subgroup of $E$. This is unavoidable, as in Remark 3.15 we show that, while the $(3,3,3)$ triangle group $G$ is not a HHG by [PS_unbounded], the direct product $\mathbb{Z}\times G$ is. In this sense, the statement of Theorem 1.4 is optimal.
@@
Focusing on quotients by the normal closure of a high power of a pseudo-Anosov element, we solve the problem for the 4-strand braid group $B_{4}$ in Theorem 4.3. The key idea is that, if a further quotient extension has hyperbolic base, then it is bounded by [neumannreeves], and hence so is the original quotient extension by Lemma 2.12. Thus the problem becomes to construct a suitable hyperbolic quotient of a five-punctured sphere, which we do using the machinery of rotating families from [dahmani:rotating] and of short HHG from [short_HHG:I].
For surfaces of higher genus, one can maybe adapt the above argument, conditionally to the residual finiteness of certain hyperbolic groups (see Remark 4.5), or try to exploit that the kernel is an increasing union of convex-cocompact subgroups (see Subsection 4.3).
@@ Acknowledgments
We would like to thank Federica Bertolotti, Roberto Frigerio, Shuhei Maruyama, Francesco Milizia, and Davide Spriano for helpful discussions. We are also grateful to Mark Hagen for explaining to us how to cubulate suitable extensions of crystallographic groups. FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund. GM is funded by an EPSRC-DTP PhD studentship.
@@ 2. Bounded extensions and quasi(homo)morphisms
is a quasimorphism of defect at most $16\delta$. Then the Busemann quasimorphism $\phi$ is the homogeneisation of $\phi_{\{x\}}$, as in Remark 2.5. One can check that $\phi$ does not depend on the choice of the sequence; moreover, an element $g\in E$ acts loxodromically on $X$ if and only if $\phi(g)\neq 0$. Note that, by construction, for every $g\in E$ one has that $|\phi(g)|\leq\mathrm{d}(x_{0},gx_{0})$; conversely, if $X$ is a quasiline, then there exists a constant $L$ such that $\mathrm{d}(x_{0},gx_{0})\leq|\mathrm{d}(gx_{0},x_{n})-\mathrm{d}(x_{0},x_{n})|+L$, for all sufficiently large $n$ (if $\{x_{0},gx_{0},x_{n}\}=\{a,b,c\}$ and $b$ is “between” $a$ and $c$, one can see this by considering the distance between $b$ and a geodesic $[a,c]$). Taking the limsup, one gets that $\mathrm{d}(x_{0},gx_{0})$ and $|\phi(g)|$ are within uniform distance.
@@
$\chi((k_{1},g_{1})\cdot(k_{2},g_{2}))=k_{1}+k_{2}+\omega(g_{1},g_{2})=\chi(k_{1},g_{1})+\chi(k_{2},g_{2})+\omega(g_{1},g_{2}).$
Thus the defect of $\chi$ is bounded by the norm of $\omega$, and so it is bounded, i.e. $\chi$ is a quasihomomorphism.
@@
We use the characterisation in the second item of Proposition 2.9. Let $\chi\colon E\to K$ be a quasihomomorphism such that $\chi|_{K}$ is the identity. Then $\chi|_{L}$ is the identity, which proves that the $L$-central extension is bounded. Composing $\chi$ with the projection $K\to K/L$, we get a quasihomomorphism $E/L\to K/L$ that is the identity on $K/L$, whcih proves that the $K/L$-central extension is bounded. ∎
Given a central extension $1\to K\to E\xrightarrow{\pi}G\to 1$, let $N\leq E$ be a normal subgroup that intersects $K$ trivially, so that $\pi\colon N\to\pi(N)$ is an isomorphism. Then there is a diagram
@@
Now $\pi\tau$ is a section for $p\colon G\to\overline{G}$, so every element of $G$ can be written uniquely as $g=\pi(n)\cdot\pi\tau(p(g))$ for some $n\in N$. We define $\sigma(g)=n\cdot\tau(p(g))$. First, note that $\sigma$ is indeed a normalised section: $\sigma(1_{G})=1_{E}$, and if $g=\pi(n)\pi\tau(p(g))$, then
@@
Moving to the second bullet, from the proof of Proposition 2.9 we see that if $\omega$ is a normalised bounded cocycle representing $E$, seen as the Cartesian product $K\times G$ with the product twisted by $\omega$, then $\chi$ can be chosen to be the projection onto the first factor, while $\pi$ is the projection onto the second factor. Now every element $g\in N$ can be written as $(b(g),\pi(g))$, and it follows that $\chi\pi^{-1}\colon\pi(N)\to K$ is just the map $\pi(g)\mapsto b(g)$. Because $N$ is a group and $\pi$ is injective, for all $g_{1},g_{2}\in N$ it holds
@@
In view of Proposition 2.16, the problem of whether the quotient central extension $\overline{E}]$ is bounded is equivalent to the problem of whether a specific quasihomomorphism $\pi(N)\to K$ extends to $G$. This problem is non-trivial. On the one hand, the easy obstruction for extendability, that is almost invariance under conjugacy by $G$, is excluded by the the first bullet of Proposition 2.16. On the other hand, the almost invariance under conjugacy is not sufficient for a quasihomomorphism to be extendable in general, as we have seen in Example 2.13. The problem of extendability has been studied at length in recent years, and produced various examples of non-extendable quasimorphisms, and criteria for extendability: see [extendable:survey] for a survey. In particular, we recall the following criteria:

### v3

@@ 1. Introduction
Central extensions $1\to K\to E\to G\to 1$ (which we often abbreviate with just $E$) are classified by their associated Euler class $\alpha_{E}\in H^{2}(G;K)$. We call a central extension bounded if its Euler class is bounded, i.e. it is represented by a bounded cocycle. One reason of interest is the fact that bounded central extensions are quasi-isometrically trivial, meaning that $E$ is quasi-isometric to the direct product $K\times G$ [Gersten_B_is_QIT]. Another reason of interest comes from the study of hierarchically hyperbolic groups, see below.
@@
By a quotient central extension we mean that there exists a diagram as follows:
@@
By Proposition 2.9 (which was recently obtained independently by Tao and Wan [tao2025properactionsfiniteproducts, Lemma 5.6]), boundedness of a central extension $E$ is equivalent to the existence of a quasihomomorphism $E\to K$ that is the identity on $K$. Quasihomomorphisms, in the sense of e.g. [fujikapo], are also relevant for quotient central extensions, and indeed we show that boundedness of a quotient central extension is equivalent to an extendability problem for a quasihomomorphism. An informal statement is as follows:
@@
We postpone to Section 4 the details of how Theorem 1.3 relates to hierarchical hyperbolicity of quotients of mapping class groups. In the other direction, under the further mild assumption that $E$ has clean containers (see Definition 3.6), we prove that a central quotient of a HHG is commensurable to a HHG:
@@
Remark 1.5.
[uniform_undistortion, Question 6.8] asks if, given a group $G$ which admits an asymptotic cone with a cut-point, then every HHG $E$ which is quasi-isometric to $\mathbb{Z}\times G$ must contain a central element $t$ such that $E/\langle t\rangle$ is virtually a HHG. Under the assumption that $E$ has clean containers, our Theorem 1.4 reduces the question to finding a central element.
Remark 1.6.
With the notation of Theorem 1.4, $G$ is a finite-index overgroup of a HHG, and therefore inherits the structure of a hierarchically hyperbolic space given by the quasi-isometric inclusion $E^{\prime}/K\hookrightarrow G$. However, even assuming that $E$ has clean containers, one cannot hope that $G$ itself is a hierarchically hyperbolic group: in Remark 3.16 we argue that the trivial central extension $E=\mathbb{Z}\times G$ of the $(3,3,3)$-triangle group $G$ is HHG, while $G$ is not by [PS_unbounded].
@@
Section 3 starts with background material on hierarchically hyperbolic groups (HHG). Roughly, a group $\Gamma$ is a HHG if there exists a collection of hyperbolic spaces $\{\mathcal{C}U\}_{U\in\mathfrak{S}}$, together with coarsely Lipschitz “coordinate projections” $\pi_{U}\colon\Gamma\to\mathcal{C}U$ satisfying various conditions. Each $U\in\mathfrak{S}$ is called a domain, and $\mathcal{C}U$ is the associated coordinate space.
@@
We uncover a connection between boundedness and hierarchical hyperbolicity of a central extension. In one direction, Theorem 3.12 shows that a bounded central extension of a HHG is itself a HHG. With an inductive argument, one can reduce to the case where the kernel is $\mathbb{Z}$, which was already settled by [uniform_undistortion, Theorem 5.14]; in turn, the proof of the latter is a refinement of [HRSS_3manifold, Corollary 4.3], which solved the case where the base is hyperbolic. We sketch here the core idea of both arguments, as it enlightens several ideas that appear repeatedly throughout our paper. From a bounded extension one gets an unbounded quasimorphism $\phi\colon E\to\mathbb{Z}$ as above, and the function $E\to\mathbb{Z}\times G$ mapping $e\in E$ to $(\phi(e),\pi(e))$ is a quasi-isometry. In other words, a bounded extension is also quasi-isometrically trivial in the sense of [Gersten_B_is_QIT], though the converse implication is not true in general (see [frigsisto, ascarimilizia] and Remark 3.14). Then the HHG structure for $E$ will have one domain for every domain of $G$, with projection to the same coordinate space factoring through the quotient map $\pi$, and one coordinate space quasi-isometric to $\mathbb{R}$ (a quasiline) to “detect” the $\mathbb{Z}$-factor. Such quasiline is built out of the quasimorphism $\phi$, using an observation of Abbott, Balasubramanya, and Osin [ABO, Lemma 4.15].
@@
In the opposite direction, Theorem 3.15 proves that, if $E$ is a HHG, then the central extension is bounded, moreover, under the further technical assumption of clean containers, the quotient is a finite-index overgroup of a HHG. To get boundedness we again look for a quasimorphism on $E$ which is the identity on $\mathbb{Z}$. Towards this, one first finds a coordinate space $\mathcal{C}U$ on which the centre acts loxodromically, which must be a quasiline; then the required quasimorphism is the Busemann quasimorphism [Manning_actions_on_hyp, Section 4.1] associated to the action of (a finite-index subgroup of) $E$ on $\mathcal{C}U$, which roughly maps each $e\in E$ to its asymptotic translation length on the quasiline.
@@
We now sketch how to prove hierarchical hyperbolicity of a finite-index subgroup of $G$. Let $\mathcal{C}U_{1},\ldots,\mathcal{C}U_{k}$ be the coordinate spaces on which $\mathbb{Z}$ acts loxodromically. By taking linear combinations of the associated Busemann quasimorphisms, and then replacing each $\mathcal{C}U_{i}$ with a new quasiline (again using [ABO, Lemma 4.15]), we can modify the HHG structure for $E$ in such a way that $\mathbb{Z}$ acts loxodromically on a single $\mathcal{C}U_{i}$, call it $\mathcal{C}U$, and with uniformly bounded orbits on every other coordinate space. Then the HHG structure of the quotient roughly coincides with what is left of the structure for $E$ after we delete $\mathcal{C}U$.
@@
In the above procedure, one has to restrict to a finite-index subgroup of $E$. More precisely, the $E$-action on $\mathfrak{S}$ permutes the collection $\mathcal{U}=\{U_{1},\ldots,U_{k}\}$, and the finite-index subgroup one considers is the kernel of the action on $\mathcal{U}$. Restricting to a subgroup is unavoidable, as in Remark 3.16 we show that, while the $(3,3,3)$ triangle group $G$ is not a HHG by [PS_unbounded], the direct product $\mathbb{Z}\times G$ is. In this sense, the statement of Theorem 1.4 is optimal.
@@
The first version of this paper contained a proof of hierarchical hyperbolicity of quotients of the braid group $B_{4}$ on four strands by a suitable power of a pseudo-Anosov (from now on, a pseudo-Anosov quotient). This fitted our framework because the mapping class group of a surface with boundary is a central extension of (a finite-index subgroup of) the mapping class group of a surface without boundaries, whose pseudo-Anosov quotients were already studied in e.g. [DGO, hhs_asdim]. The key idea was that, if a further quotient extension has hyperbolic base, then it is bounded by [neumannreeves], and therefore the original quotient extension is bounded by Lemma 2.12. Thus the problem reduced to construct a suitable hyperbolic quotient of the base of the extension, which involved the machinery of rotating families from [dahmani:rotating] and of short HHG from [short_HHG:I]. This approach could possibly be adapted to higher-genus surfaces with boundary, conditionally to the residual finiteness of certain hyperbolic groups (see Remark 4.4).
Since then, Tao proved that pseudo-Anosov quotients of any mapping class group are hierarchically hyperbolic [tao2025extensiontheoremquasimorphisms, Corollary 1.7]. Tao’s argument uses our Proposition 1.2 and Theorem 1.3 to reduce the problem to the extendability of a certain quasimorphism; then he proves a theorem which unifies most known extendability results for quasimorphisms. We still decided to include a streamline of our original arguments as a proof of concept, see Theorem 4.3.
@@ Acknowledgments
We would like to thank Federica Bertolotti, Roberto Frigerio, Shuhei Maruyama, Francesco Milizia, Davide Spriano, and Bingxue Tao for helpful discussions. We are also grateful to Mark Hagen for explaining to us how to cubulate suitable extensions of crystallographic groups. FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund. GM is funded by an EPSRC-DTP PhD studentship.
@@ 2. Bounded extensions and quasi(homo)morphisms
is a quasimorphism of defect at most $16\delta$. Then the Busemann quasimorphism $\phi$ is the homogeneisation of $\phi_{\{x\}}$, as in Remark 2.5. One can check that $\phi$ does not depend on the choice of the sequence; moreover, an element $g\in E$ acts loxodromically on $X$ if and only if $\phi(g)\neq 0$. Note that, by construction, for every $g\in E$ one has that $|\phi(g)|\leq\mathrm{d}(x_{0},gx_{0})$; conversely, if $X$ is a quasiline (i.e. if $X$ is quasi-isometric to $\mathbb{R}$), then there exists a constant $L$ such that $\mathrm{d}(x_{0},gx_{0})\leq|\mathrm{d}(gx_{0},x_{n})-\mathrm{d}(x_{0},x_{n})|+L$, for all sufficiently large $n$ (if $\{x_{0},gx_{0},x_{n}\}=\{a,b,c\}$ and $b$ is “between” $a$ and $c$, one can see this by considering the distance between $b$ and a geodesic $[a,c]$). Taking the limsup, one gets that $\mathrm{d}(x_{0},gx_{0})$ and $|\phi(g)|$ are within uniform distance.
@@
$\psi((k_{1},g_{1})\cdot(k_{2},g_{2}))=k_{1}+k_{2}+\omega(g_{1},g_{2})=\psi(k_{1},g_{1})+\psi(k_{2},g_{2})+\omega(g_{1},g_{2}).$
Thus the defect of $\psi$ is bounded by the norm of $\omega$, and so it is bounded, i.e. $\psi$ is a quasihomomorphism.
@@
We use the characterisation in the second item of Proposition 2.9. Let $\chi\colon E\to K$ be a quasihomomorphism such that $\chi|_{K}$ is the identity. Then $\chi|_{L}$ is the identity, which proves that the $L$-central extension is bounded. Composing $\chi$ with the projection $K\to K/L$, we get a quasihomomorphism $E/L\to K/L$ that is the identity on $K/L$, which proves that the $K/L$-central extension is bounded. ∎
2.1. (Bounded) quotient extensions
Let $1\to K\to E\xrightarrow{\pi}G\to 1$ be a central extension, and let $N\leq E$ be a normal subgroup that intersects $K$ trivially, so that $\pi\colon N\to\pi(N)$ is an isomorphism. Then there is a diagram
@@
Now $\pi\tau$ is a section for $p\colon G\to\overline{G}$, so every element of $G$ can be written uniquely as $g=\pi(n)\cdot\pi\tau(p(g))$ for some $n\in N$. We define $\sigma(g)=n\cdot\tau(p(g))$, so $\sigma\colon G\to E$. First, note that $\sigma$ is indeed a normalised section: $\sigma(1_{G})=1_{E}$, and if $g=\pi(n)\pi\tau(p(g))$, then
@@
Moving to the second bullet, from the proof of Proposition 2.9 we see that if $\omega$ is a normalised bounded cocycle representing $E$, seen as the Cartesian product $K\times G$ with the product twisted by $\omega$, then $\chi$ can be chosen to be the projection onto the first factor, while $\pi$ is the projection onto the second factor. Now every element $g\in N$ can be written uniquely as $(b(g),\pi(g))$, for some $b(g)\in K$. It follows that $\chi\pi^{-1}\colon\pi(N)\to K$ is just the map $\pi(g)\mapsto b(g)$. Because $N$ is a group and $\pi$ is injective, for all $g_{1},g_{2}\in N$ it holds
@@
2.2. Extending quasimorphisms
In view of Proposition 2.16, the problem of whether the quotient central extension $\overline{E}]$ is bounded is equivalent to the problem of whether a specific quasihomomorphism $\pi(N)\to K$ extends to $G$. This problem is non-trivial. On the one hand, the easy obstruction for extendability, that is almost invariance under conjugacy by $G$, is excluded by the the first bullet of Proposition 2.16. On the other hand, almost invariance under conjugacy is not sufficient for a quasihomomorphism to be extendable in general, as we have seen in Example 2.13.
Remark 2.18.
In the context of Proposition 2.16, the quasihomomorphism one has to extend is always such that its boundary is extendable. This shows a stark difference between the problem of extending bounded cohomology classes and quasihomomorphisms [extendable:survey, end of Section 1.3].
2.2.1. Extendability results in the literature
The problem of extending quasihomomorphisms has been studied at length in recent years, and produced various examples of non-extendable quasimorphisms, and criteria for extendability, see [extendable:survey] for a survey. In particular, we recall the following criteria:
@@ 2.2.1. Extendability results in the literature
Recently, Tao developed a very general framework that unifies the results for hyperbolically embedded subgroups and for normal subgroups, and has many new consequences [tao2025properactionsfiniteproducts].
2.2.2. Extending quasimorphisms from a union
We conclude the section by proving a new extendability result, which applies to subgroups obtained as an increasing union of subgroups where quasimorphisms can be extended. We hope that this will find applications towards extending quasimorphisms from “small-cancellation-like” normal subgroups.
Lemma 2.21.
Proof.
By using homogeneous representatives for one direction, and integer parts for the other, as we did before, we can replace $\mathbb{Z}$ by $\mathbb{R}$ in the statement, thus reducing to prove the following. Let $\chi\colon N\to\mathbb{R}$ be a homogeneous quasimorphism. Suppose that $N$ can be expressed as a directed union of subgroup $\{N_{i}\}_{i\in I}$, and that for each $i\in I$ there is a homogeneous quasimorphism $X_{i}\colon G\to\mathbb{R}$ that extends $\chi|_{N_{i}}$, such that the defect of the $X_{i}$ is uniformly bounded. Then $\chi$ extends to a homogeneous quasimorphism $X\colon G\to\mathbb{R}$. Moreover, because $G$ is finitely generated, thus countable, we can assume that the directed set $I$ is just $\mathbb{N}$ with its well order.
$\alpha(g^{p})=\alpha\left(\prod\limits_{j=1}^{n}x_{j}^{p_{j}}\right)$